OpenStudy (amonoconnor):

For the brave at heart... I need help checking the derivative(s) of the following function: "y = x/(x^2 - 4)" The following are what I got for the first & second derivatives, respectively: "y' = (-x^2 - 4)/(x^2 - 4)^2" , "y" = [(2(x^2 - 4))(x(x^2 - 4) - x(-x^2 - 4))]/(x^2 - 4)^2 Yeah... Disgusting, right? That's why I'm hoping to have some burial land soul go through this one and check my work, because I know even a tiny mistake can ruin a full-page problem. (To be continued... )

OpenStudy (amonoconnor):

*Note: 1. I used the Quotient Rule to find the derivative I did. 2. No, what I have written is not the initial second derivative- I did as much combining like terms/ factoring as I knew how to. The 'original' second derivative, without moving, adding, or changing anything from the format of the Q.R. equation is: "[-2x((x^2 - 4)^2) - (-x^2 - 4)(2(x^2 - 4))(2x)]/(x^2 - 4)^4"

OpenStudy (amonoconnor):

Any and all help is greatly appreciated! :)

OpenStudy (triciaal):

y = x/(x^2 -4) let x =u let (x^2-4)= v dy/dx = (v*du/dx-u*dv/dx) / v^2 du/dx = 1 dv/dx = 2x v^2 = (x^4 - 8x^2+ 16) y' = (x^2-4)(1)-x(2x) all divided by (x^4-8x^2+16) (x^2 -4 -2x^2)/(x^4-8x^2+16) (-3x^2 -4)/(x^4 -8x^2+16)

OpenStudy (amonoconnor):

?? So... Is my work right? :) Oh, and I think that it should be -1x^2 in the last step you did. Right? 1x^2 - 2x^2 = -1x^2 ?

OpenStudy (anonymous):

Use wolframalpha to check your answers http://www.wolframalpha.com/input/?i=++derivative+x%2F%28x^2+-+4%29

OpenStudy (amonoconnor):

Ohhhh!! Awesome! This website is amazing!! :) Thanks @eliassaab !

OpenStudy (anonymous):

Yw

OpenStudy (anonymous):

You can also practice calculus on my interactive site http://www.saab.org/calculus.cgi

OpenStudy (amonoconnor):

@wio any thoughts? :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=d%2Fdx++x%2F(x%5E2+-+4) This is what I think.

OpenStudy (amonoconnor):

Touché. ;) Alright, guess I'll go that route.